# Three meta-questions on infinite-domain Constraint Satisfaction Problems

### v1

@@ 4.3. Putting everything together
Regarding item 2 (homogeneity and the Ramsey property), one first observes that, since $(\mathbb{Q};<)^{\mathfrak{B}}_{I_{4}}$ is an expansion of $\mathfrak{B}^{\prime}$ by two relations that are preserved by all embeddings, we can ignore these relations and only prove the statement for $\mathfrak{B}^{\prime}$. Then homogeneity follows directly from Proposition 4.5(4) because $(\mathbb{Q};<)$ is homogeneous. In the second part, the backward direction follows directly from Proposition 4.5(6) because $(\mathbb{Q};<)$ is homogeneous Ramsey [54]. The forward direction can be proved by hand, by instantiating the Ramsey property for the age of $\mathfrak{B}^{\prime}$ on structures where $E$ interprets as the diagonal relation.
@@
Regarding item 3 (homogeneity and finite boundedness), homogeneity follows directly from the previous item. As for finite boundedness, it is a consequence of Proposition 4.5(5) (which immediately yields finite boundedness of $\mathfrak{B}^{\prime}$), and since we take an expansion by relations definable by a Boolean combination of equality atoms. More precisely, recall that up to isomorphism, the finite substructures of a finitely bounded structure are precisely the finite models of a universal sentence. By adding new clauses defining the relations $\neq$ and $I_{4}$ in terms of equalities to the universal sentence describing the finite substructures of $\mathfrak{B}^{\prime}$, we obtain a universal sentence describing the finite substructures of $(\mathbb{Q};<)^{\mathfrak{B}}_{I_{4}}$.
@@
Regarding item 4 (pp-construction of $\mathfrak{K}_{3}$), we use a characterization via the pseudo-Siggers identity from [7, Theorems 1.3 and 3.4] and the fact that the structure $(\mathbb{Q};\neq,I_{4})$ is an $\omega$-categorical model-complete core that does not pp-construct $\mathfrak{K}_{3}$ [16, Theorems 12.0.1, 12.7.3, 12.9.2, and Corollary 6.4.4]. One direction is simple (and does not use the assumption on the orbit growth), because $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ clearly pp-constructs $\mathfrak{A}$; the pp-construction is given by $\mathfrak{A}=(\smash{\mathbb{Q}^{\mathfrak{A}}/E^{\smash{\mathbb{Q}^{\mathfrak{A}}_{I_{4}}}})|^{\tau}}$ (since taking factors and reducts are pp-constructions [9]). Thus, if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$, then so does $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$. For the other direction, one assumes that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$, in which case Theorem 1.3 in [7] yields polymorphisms in its model-complete core satisfying the pseudo-Siggers identity. These are then composed with polymorphisms of $(\mathbb{Q};\neq,I_{4})$ witnessing the pseudo-Siggers identity, which ultimately yields the satisfaction of the pseudo-Siggers identity by polymorphisms of the model-complete core of $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$. The proof of Proposition 4.5(3) shows that also the structure $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ has less than double exponential orbit growth. It then follows from Theorem 3.4 in [7] (whose item (ii) is equivalent to the pp-construction of $\mathfrak{K}_{3}$ by Theorem 1.8 in [9]) that $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. ∎
We conclude this section with two examples.
@@
Consider $\mathfrak{A}_{1}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(\{0,1\};\{1\},\{(x,y,z)\in\{0,1\}^{3}\mid x+y+z=0\bmod 2\})$. This structure is a finitely bounded homogeneous model-complete core with algebraicity (since we do not exclude finite structures from these definitions); its CSP is solvable using Gaussian elimination over $\mathbb{F}_{2}$. Note that $\mathfrak{A}_{1}$ is preserved by the ternary symmetric operation $(x,y,z)\mapsto x+y+z\bmod 2$, which means that its polymorphism clone satisfies the h1-identities $f(x,y,z)\approx f(y,z,x)\approx f(z,x,y)$. Since pp-constructions preserve the satisfaction of h1-identities in polymorphism clones [9] and $\neq$ over an infinite set is not preserved by any operation $f$ satisfying these identities, $\mathfrak{A}_{1}$ does not pp-construct $\mathbb{Q}^{\mathfrak{A}_{1}}$. Hence, our Datalog-reduction is not subsumed by pp-constructions.
Next, consider $\mathfrak{A}_{2}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(\mathbb{Q};<,\{(x,y,z)\in\mathbb{Q}^{3}\mid x\geq y\text{ or }x\geq z\})$. This structure is a finitely bounded homogeneous model-complete core without algebraicity; its CSP is definable in fixed-point logic [34, Proposition 3.7]. Note that $\mathfrak{A}_{2}$ is preserved by the binary commutative operation $(x,y)\mapsto\min(x,y)$, and hence its polymorphism clone satisfies the identity $f(x,y)\approx f(y,x)$. For the same reason as above, $\mathfrak{A}_{2}$ does not pp-construct $\mathbb{Q}^{\mathfrak{A}_{2}}$.
5. $\omega$-categorical sandwiches for PCSPs
@@ 5.1. PCSP polymorphisms and full powers
We extend the notion of a polymorphism to PCSP templates $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ in the usual way: a $k$-ary polymorphism of $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ is a homomorphism from a finite power of $\mathfrak{S}_{1}$ to $\mathfrak{S}_{2}$; we denote by $\mathrm{Pol}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ the set of all such polymorphisms. Also the satisfaction of height-1 conditions can be generalized to the PCSP setting in a straightforward manner. In the case of nested identities, we have to be more careful because the composition of polymorphisms between different structures is not defined. We say that a height-1 condition $\Sigma$ is satisfied in $\mathrm{Pol}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ modulo a set $H$ of unary operations on $S_{2}$ if the function symbols in $\Sigma$ can be interpreted as elements of $\mathrm{Pol}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ such that for each identity $s\approx t$ in $\Sigma$ there exist $\alpha,\beta\in H$ such that $\alpha\circ s=\beta\circ t$ holds for any evaluation of variables in $S_{1}$.
Our basic tool for creating $\omega$-categorical sandwiches for PCSPs that are not finitely tractable are full powers (see, e.g., Bodirsky [16, Section 3.5]). Roughly speaking, a full power ${\mathfrak{A}}^{[d]}$ is a higher-dimensional representation of $\mathfrak{A}$ that can be pp-constructed from it and vice versa; its central feature is that its polymorphisms are the polymorphisms of $\mathfrak{A}$ acting on $d$-tuples, thus allowing to factor it by the orbit-equivalence of $\mathrm{Aut}(\mathfrak{A})$ on $d$-tuples (rather than on elements). In other words, the orbits of $d$-tuples of $\mathfrak{A}$ are represented by orbits of $1$-tuples of ${\mathfrak{A}}^{[d]}$, which is crucial in applications of the sandwiching method.
Let $\mathfrak{A}$ be a relational structure with a signature $\tau$ and let $d\in\mathbb{N}$ be arbitrary. The $d$-th full power of $\mathfrak{A}$, denoted ${\mathfrak{A}}^{[d]}$, is the structure with domain $A^{d}$ and the following relations for every $k\in[d]$:
@@
    $R^{{\mathfrak{A}}^{[d]}}_{\iota}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\{(a_{1}\dots,a_{d})\in A^{d}\mid(a_{\iota(1)},\dots,a_{\iota(k)})\in R^{\mathfrak{A}}\};\hskip 99.58464pt$
@@
    $\widehat{R}^{{\mathfrak{A}}^{[d]}}_{\iota}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\{((a^{1}_{1},\dots,a^{1}_{m}),\dots,(a^{k}_{1},\dots,a^{k}_{d}))\in(A^{d})^{k}\mid(a^{1}_{\iota(1)},\dots,a^{k}_{\iota(k)})\in R^{\mathfrak{A}}\};$
  • for all functions $\iota,\iota^{\prime}\colon[k]\rightarrow[d]$, define the binary compatibility relation
    $S^{{\mathfrak{A}}^{[d]}}_{\iota,\iota^{\prime}}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\{(a^{1}_{1}\dots,a^{1}_{d}),(a^{2}_{1},\dots,a^{2}_{d})\in(A^{d})^{2}\mid\forall i\in[k]\colon a^{1}_{\iota(i)}=a^{2}_{\iota^{\prime}(i)}\}.\hskip 35.56593pt$
Below we give several crucial properties of full powers.
@@
    ${\mathfrak{A}}^{[d]}$ is pp-constructible from $\mathfrak{A}$ and vice versa.
@@ 5.2. Our sandwich recipe
Let $\mathfrak{A}$ be a reduct of a linearly ordered finitely bounded homogeneous structure $\mathfrak{B}$ such that the polymorphisms of $\mathfrak{A}$ preserve $\neq$. Let $d\in\mathbb{N}$ be such that the bounds of $\mathfrak{B}$ have size $\leq d$ and its relations are of arity $\leq d-1$. Then there exists a finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ such that:
@@
    ${\mathfrak{A}}^{[d]}$ is a sandwich for $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$;
@@
    Every finite sandwich for $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ pp-constructs $\mathfrak{K}_{3}$.
@@
Clearly, for every finite substructure $\mathfrak{S}$ of $\mathfrak{A}$, we have ${\mathfrak{S}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. We show that there exists a finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ satisfying the second item. Since $\mathfrak{A}$ is $\omega$-categorical and its polymorphisms preserve $\neq$, by a standard compactness argument there exists a finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ such that $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ preserves $\neq$. Moreover, we can without loss of generality assume $|S|\geq 3$. Now suppose that there is a finite sandwich $\mathfrak{D}$ of $({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ that does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 3.1, there exists a cyclic operation $f$ in $\mathrm{Pol}(\mathfrak{D})$ of some arity $n$. Take any homomorphisms $g\colon{\mathfrak{S}}^{[d]}\rightarrow\mathfrak{D}$ and $h\colon\mathfrak{D}\rightarrow\smash{{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}$. Then, by setting $f^{\prime}(x_{1},\dots,x_{n})\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=h\circ f(g(x_{1}),\dots,g(x_{n})),$ we get a cyclic $f^{\prime}\in\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ of arity $n$. One can show that every polymorphism from ${\mathfrak{S}}^{[d]}$ to ${\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ can be lifted to a homomorphism from $\mathfrak{S}$ to $\mathfrak{A}$ and that any identity that is satisfied in $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ is also satisfied in $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ modulo $\mathrm{Aut}(\mathfrak{B})$. Hence, there is some $n$-ary $g\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that is pseudo-cyclic modulo $\mathrm{Aut}(\mathfrak{B})$. One easily sees that any operation that is pseudo-cyclic modulo some order-preserving automorphism group must be cyclic. But any cyclic operation on a set of size at least 3 does not preserve $\neq$, a contradiction. ∎
The following statement is folklore; see, e.g., [16, Section 2.3.6].

### v2

@@ 4.3. Removing algebraicity
Regarding item 2 (homogeneity and the Ramsey property), one first observes that, since $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}}$ is an expansion of $\mathfrak{B}^{\prime}$ by two relations that are preserved by all embeddings, we can ignore these relations and only prove the statement for $\mathfrak{B}^{\prime}$. Then homogeneity follows directly from Proposition 4.5(4) because $(\mathbb{Q};<)$ is homogeneous. In the second part, the backward direction follows directly from Proposition 4.5(6) because $(\mathbb{Q};<)$ is homogeneous Ramsey [52]. The forward direction can be proved by hand, by instantiating the Ramsey property for the finite substructures of $\mathfrak{B}^{\prime}$ where $E$ interprets as the diagonal relation.
@@
Regarding item 3 (finite boundedness), this is a consequence of Proposition 4.5(5) (which immediately yields finite boundedness of $\mathfrak{B}^{\prime}$) and the fact that we take an expansion by relations definable by a Boolean combination of equality atoms. More precisely, recall that up to isomorphism, the finite substructures of a finitely bounded structure are precisely the finite models of a universal sentence. By adding new clauses defining the relations $\neq$ and $I_{4}$ in terms of equalities to the universal sentence describing the finite substructures of $\mathfrak{B}^{\prime}$, we obtain a universal sentence describing the finite substructures of $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}}$.
@@
Regarding item 4 (pp-construction of $\mathfrak{K}_{3}$), we use a characterization via the pseudo-Siggers identity from [7, Theorems 1.3 and 3.4] and the fact that the structure $(\mathbb{Q};\neq,I_{4})$ is an $\omega$-categorical model-complete core that does not pp-construct $\mathfrak{K}_{3}$ [15, Theorems 12.0.1, 12.7.3, 12.9.2, and Corollary 6.4.4]. One direction is simple (and does not use the assumption on the orbit growth), because $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ clearly pp-constructs $\mathfrak{A}$; the pp-construction is given by $\mathfrak{A}=(\smash{(\mathbb{Q}\wr\mathfrak{A})_{\neq}/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}})|^{\tau}}$ (since taking factors and reducts are pp-constructions [9]). Thus, if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$, then so does $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. For the other direction, one assumes that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$, in which case Theorem 1.3 in [7] yields polymorphisms in its model-complete core satisfying the pseudo-Siggers identity. These are then composed with polymorphisms of $(\mathbb{Q};\neq,I_{4})$ witnessing the pseudo-Siggers identity, which ultimately yields the satisfaction of the pseudo-Siggers identity by polymorphisms of the model-complete core of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. The proof of Proposition 4.5(3) shows that also the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ has less than double exponential orbit growth. It then follows from Theorem 3.4 in [7] (whose item (ii) is equivalent to the pp-construction of $\mathfrak{K}_{3}$ by Theorem 1.8 in [9]) that $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. ∎
We conclude this section with an example.
@@
Consider the two structures
$\mathfrak{A}_{1}\coloneqq(\{0,1\};\{1\},\{(x,y,z)\in\{0,1\}^{3}\mid x+y+z=0\bmod 2\}),$
$\mathfrak{A}_{2}\coloneqq(\mathbb{Q};<,\{(x,y,z)\in\mathbb{Q}^{3}\mid x\geq y\text{ or }x\geq z\}).$
The former is a finitely bounded homogeneous model-complete core with algebraicity (due to it being finite) and the latter is a finitely bounded homogeneous model-complete core without algebraicity. Note that $\mathfrak{A}_{1}$ is preserved by the symmetric operation $(x_{1},\dots,x_{n})\mapsto x_{1}+\cdots+x_{n}\bmod 2$ for every odd $n\geq 2$ and $\mathfrak{A}_{2}$ is preserved by the symmetric operation $(x_{1},\dots,x_{n})\mapsto\min\{x_{1},\dots,x_{n}\}$ for every $n\geq 2$. Thus, in both cases the polymorphism clone satisfies the cyclic identity of arity $n$ for some $n\geq 2$. Since pp-constructions preserve the satisfaction of height-1 identities in polymorphism clones [9] and no cyclic operation $f$ of arity $n\geq 2$ preserves $\neq$ over an infinite set, $\mathfrak{A}_{i}$ does not pp-construct $(\mathbb{Q}\wr\mathfrak{A}_{i})_{\neq}$ for both $i\in[2]$. Hence, our Datalog-reduction in Theorem 2.2 is not subsumed by gadget reductions (pp-constructions).
5. $\omega$-categorical cheeses for PCSPs
@@ 5.1. PCSP polymorphisms and full powers
We extend the notion of a polymorphism to PCSP templates $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ in the usual way: a polymorphism of $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ is a homomorphism from a finite power of $\mathfrak{S}_{1}$ to $\mathfrak{S}_{2}$; we denote by $\mathrm{Pol}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ the set (minion) of all such polymorphisms. Also the satisfaction of height-1 conditions can be generalized to the PCSP setting in a straightforward manner. In the case of nested identities, one has to be more careful because the composition of polymorphisms between different structures is not defined.We say that a height-1 condition $\Sigma$ is satisfied in $\mathrm{Pol}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ modulo a set $H$ of unary operations on $S_{2}$ if the function symbols in $\Sigma$ can be interpreted as elements of $\mathrm{Pol}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ such that for each identity $s\approx t$ in $\Sigma$ there exist $\alpha,\beta\in H$ such that $\alpha\circ s=\beta\circ t$ holds for any evaluation of variables in $S_{1}$.
Our basic tool for creating $\omega$-categorical cheeses for PCSPs that are not finitely tractable are full powers (see, e.g., Bodirsky [15, Section 3.5]). Roughly speaking, a full power ${\mathfrak{A}}^{[d]}$ is a higher-dimensional representation of $\mathfrak{A}$ that can be pp-constructed from it and vice versa; its central feature is that its polymorphisms are the polymorphisms of $\mathfrak{A}$ acting on $d$-tuples, thus allowing to factor it by the orbit-equivalence of $\mathrm{Aut}(\mathfrak{A})$ on $d$-tuples (rather than on elements). In other words, the orbits of $d$-tuples of $\mathfrak{A}$ are represented by orbits of $1$-tuples of ${\mathfrak{A}}^{[d]}$, which is crucial in applications of the sandwich method. Let $\mathfrak{A}$ be a relational structure with a signature $\tau$ and let $d\in\mathbb{N}$ be arbitrary. The $d$-th full power of $\mathfrak{A}$, denoted ${\mathfrak{A}}^{[d]}$, is the structure with domain $A^{d}$ and the following relations for every $k\in[d]$:
@@
    $R^{{\mathfrak{A}}^{[d]}}_{\iota}\coloneqq\{(a_{1}\dots,a_{d})\in A^{d}\mid(a_{\iota(1)},\dots,a_{\iota(k)})\in R^{\mathfrak{A}}\};$
@@
    $\widehat{R}^{{\mathfrak{A}}^{[d]}}_{\iota}\coloneqq\{((a^{1}_{1},\dots,a^{1}_{d}),\dots,(a^{k}_{1},\dots,a^{k}_{d}))\in(A^{d})^{k}\mid(a^{1}_{\iota(1)},\dots,a^{k}_{\iota(k)})\in R^{\mathfrak{A}}\};$
  • for all functions $\iota,\iota^{\prime}\colon[k]\rightarrow[d]$, the binary compatibility relation
    $S^{{\mathfrak{A}}^{[d]}}_{\iota,\iota^{\prime}}\coloneqq\{((a^{1}_{1}\dots,a^{1}_{d}),(a^{2}_{1},\dots,a^{2}_{d}))\in(A^{d})^{2}\mid\forall i\in[k]\colon a^{1}_{\iota(i)}=a^{2}_{\iota^{\prime}(i)}\}.$
Below we give several important properties of full powers.
@@
    ${\mathfrak{A}}^{[d]}$ is (factorisation-free) pp-interpretable from $\mathfrak{A}$ and vice versa.
@@ 5.2. Our sandwich recipe
Let $\mathfrak{A}$ be a reduct of a linearly ordered finitely bounded homogeneous structure $\mathfrak{B}$ such that one of the relations of $\mathfrak{A}$ interprets as $\neq$. Let $d\in\mathbb{N}$ be such that the bounds of $\mathfrak{B}$ have size $\leq d$ and its relations are of arity $\leq d-1$. Then, for every finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ with $|S|\geq 3$:
@@
    ${\mathfrak{A}}^{[d]}$ is a cheese for $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$;
@@
    Every finite cheese for $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ pp-constructs $\mathfrak{K}_{3}$.
@@
Clearly, for every finite substructure $\mathfrak{S}$ of $\mathfrak{A}$, we have $\smash{{\mathfrak{S}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}$. Now suppose that there is a finite cheese $\mathfrak{D}$ of $({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ that does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 3.1, there exists a cyclic operation $f$ in $\mathrm{Pol}(\mathfrak{D})$ of some arity $n$. Take any homomorphisms $g\colon{\mathfrak{S}}^{[d]}\rightarrow\mathfrak{D}$ and $h\colon\mathfrak{D}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Then, by setting $f^{\prime}(x_{1},\dots,x_{n})\coloneqq h\circ f(g(x_{1}),\dots,g(x_{n})),$ we get a cyclic $f^{\prime}\in\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ of arity $n$. One can show that every polymorphism from ${\mathfrak{S}}^{[d]}$ to ${\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ can be lifted to a polymorphism from $\mathfrak{S}$ to $\mathfrak{A}$ and that any height-1 identity that is satisfied in $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ is also satisfied in $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ modulo $\mathrm{Aut}(\mathfrak{B})$. Hence, there is some $n$-ary $g\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that is pseudo-cyclic modulo $\mathrm{Aut}(\mathfrak{B})$. One easily sees that any operation that is pseudo-cyclic modulo some order-preserving automorphism group must be cyclic. This means that we have successfully lifted the cyclic identity from $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ to $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$. But any cyclic operation on a set of size at least 3 does not preserve $\neq$, a contradiction. ∎
The following statement is folklore; see e.g. [15, Section 2.3.6].